OpenStudy (anonymous):

In 4 years time a mother will be three times as old as her son. Fours years ago she was five times as old as her son. Find their present ages.

Parth (parthkohli):

Let's say that the mother is \(M\), and the son is \(S\).\[M + 4 = 3(S + 4) \]\[M - 4= 5(S - 4) \]Can you solve?

OpenStudy (anonymous):

Yes! OMG THANK YOUUU @ParthKohli

Parth (parthkohli):

You're welcome :)